The centre of a wheel rolling on a plane surface moves with a speed v0
. A particle on the rim of the wheel
at the same level as the centre will be moving at speed

Grade:

1 Answers

Prity
6 years ago
As it is on the rim of the wheel so the distance means r =r^2+r^2. And here we get r become root2r and from equation we know v=w*r we get the answer v=root 2u